OpenStudy (scarlettthecamel):

Factorise completely x^2+x-12

DivineSolar (divinesolar):

Factoring x2+x-12 The first term is, x2 its coefficient is 1 . The middle term is, +x its coefficient is 1 . The last term, "the constant", is -12 Step-1 : Multiply the coefficient of the first term by the constant 1 • -12 = -12 Step-2 : Find two factors of -12 whose sum equals the coefficient of the middle term, which is 1 . Try doing step 2.

OpenStudy (scarlettthecamel):

Okay.

DivineSolar (divinesolar):

I'll do 2 and you do the other 2, -12 + 1 = -11 -6 + 2 = -4

OpenStudy (scarlettthecamel):

I don't understand whats goingon lol

DivineSolar (divinesolar):

Ok fine here, Step-2 : Find two factors of -12 whose sum equals the coefficient of the middle term, which is 1 . -12 + 1 = -11 -6 + 2 = -4 -4 + 3 = -1 -3 + 4 = 1 That's it

DivineSolar (divinesolar):

However you have to do this, Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -3 and 4

OpenStudy (scarlettthecamel):

Is there a way u can explain it without the mathematical terms becoz it confuses me lol

DivineSolar (divinesolar):

Not really. I like to explain it mathematically it is my part you know?

DivineSolar (divinesolar):

Try this. Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -3 and 4 x2 - 3x + 4x - 12 Step-4 : Add up the first 2 terms, pulling out like factors : x • (x-3) Add up the last 2 terms, pulling out common factors : 4 • (x-3) Step-5 : Add up the four terms of step 4 :

OpenStudy (scarlettthecamel):

(2x-3)(4x-3) ?????????

DivineSolar (divinesolar):

No it would be (x + 4) (x - 3)

DivineSolar (divinesolar):

(x + 4) * (x - 3)

DivineSolar (divinesolar):

See how i got that?

OpenStudy (scarlettthecamel):

not really

DivineSolar (divinesolar):

Well look over the steps :).

OpenStudy (scarlettthecamel):

k

DivineSolar (divinesolar):

Well i need to help someone else, hope i helped you out!